Wat do you think the maximum area of a rectangle that has a perimeter of 16??
Found 2 solutions by swincher4391, ewatrrr:
Answer by swincher4391(1107)

The maximum area of a rectangle is actually achieved when that rectangle is a square.
P = 4s
16 = 4s
s = 4
A = 4^2 = 16

Answer by ewatrrr(24785)

 
Hi,
maximum area of a rectangle that has a perimeter of 16
Square with each side = 4
4*4 = 16unit^2, maximum Area
Note:
P = 2*L + 2*w = 16
L + w = 8
Other areas for example:
3*5 = 15
2*6 = 12